A dairy keeps two kinds of milk on hand, skim milk that has 0.3% butterfat and whole milk that has 3.8% butter fat. How many gallons of each type of milk does a company need to produce 380 gallons of 1% milk for a grocery store?
Answer by addingup(3677)

0.003x+0.038(380-x) = 0.01(380)
0.003x+14.44-0.038x = 3.8
-0.035x = -10.64
x = 304
304 gallons of 0.3% with 76 gallons of 3.8% to get 380 gall at 1%